# Slant sums of quiver gauge theories

### v1

@@ 3.5 From branching to factorization
Due to form of the obstruction theory, the $q=\hbar$ specialization of vertex functions with descendant $1$ depends only on $\hbar$. In particular, it does not depend on any framing parameters.
@@
Another interesting special case occurs when $\mathsf{w}^{(2)}_{\star_{2}}=1$ and $\mathsf{w}^{(2)}_{i}=0$ otherwise, which most closely resembles the original slant sum constructions of [38].
@@ 4.1 Factorization in zero-dimensional case
In this section, we will follow standard notation and denote simple roots by $\alpha_{i}$ and fundamental weights by $\varpi_{i}$. We will also denote $\kappa=q/\hbar$.
Corollary 3.5 provides a strategy to approach Conjecture 1.3. This conjecture is known only for $A$ and $D$ type quiver varieties with framings only at minuscule vertices, for which it was proven in [14] and [17] respectively.
@@
$\mathsf{V}_{\mathcal{M}}$ $=\sum_{d_{i,j}}\left(\frac{q}{\hbar}\right)^{N(d)}z_{1}^{d_{1,1}}z_{2}^{d_{2,1}+d_{2,2}}z_{3}^{d_{3,1}}\frac{\left(\hbar\right)_{d_{2,1}}\left(\hbar^{2}\right)_{d_{2,2}}}{\left(q\right)_{d_{2,1}}\left(q\hbar\right)_{d_{2,2}}}\cdot$
@@
$\mathsf{V}_{\mathcal{M}}=\frac{\Phi\left(\hbar\kappa z_{2}\right)\Phi(\hbar z_{1}z_{2})\Phi(\hbar\kappa^{2}z_{2}z_{3})\Phi(\hbar\kappa z_{1}z_{2}z_{3})}{\Phi(\kappa z_{2})\Phi(z_{1}z_{2})\Phi(\kappa^{2}z_{2}z_{3})\Phi(\kappa z_{1}z_{2}z_{3})}$
The weights $\lambda$ and $\mu$ in this case are $\lambda=\varpi_{2}$ and $\mu=-\varpi_{2}$. It is easy to see that
$(\alpha,\mu)=\begin{cases}-1&\text{if $\alpha\in\{\alpha_{2},\alpha_{1}+\alpha_{2},\alpha_{2}+\alpha_{3},\alpha_{1}+\alpha_{2}+\alpha_{3}\}$}\\ \geq 0&\text{otherwise}\end{cases}$
for any positive root $\alpha$. Recalling that $e^{\alpha_{1}}=\kappa^{-1}z_{1}$, $e^{\alpha_{2}}=\kappa z_{2}$, and $e^{\alpha_{3}}=\kappa z_{3}$, this confirms Conjecture 1.3 in this case.
Suppose that $\mathcal{M}^{(1)}$ and $\mathcal{M}^{(2)}$ are zero-dimensional quiver varieties whose vertex functions are known to factorize into a product of $q$-binomial functions. By the discussion above, they could be type $A$ or $D$ quiver varieties with framings only at minuscule vertices. Suppose that $\star_{1}$ and $\star_{2}$ are compatible vertices, that $\mathcal{M}^{(1)}$ is split over $\star_{1}$, and that $\theta^{(2)}=\pm(1,1,\ldots,1)$. Consider $\mathcal{M}:=\mathcal{M}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathcal{M}^{(2)}$.
With these assumptions, Corollary 3.5 implies the following.
@@
Let $\mathcal{M}^{(1)}$ be the type $A_{3}$ quiver variety with $\mathsf{v}^{(1)}=(1,2,1)$ and $\mathsf{w}^{(1)}=(0,1,0)$. Let $\mathcal{M}^{(2)}$ be the type $A_{1}$ quiver variety with $\mathsf{v}^{(2)}=(2)$ and $\mathsf{w}^{(2)}=(2)$. Let $\star_{1}$ be the second vertex and let $\star_{2}$ be the first (and only) vertex. As a $\mathsf{T}^{(1)}$-representation, $\mathcal{V}_{\star_{1}}=1+\hbar$. So $\mathcal{M}^{(1)}$ is split over $\star_{1}$.
Let $\mathcal{M}=\mathcal{M}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathcal{M}^{(2)}$. It is a $D_{4}$ quiver variety with $\mathsf{v}=(1,2,1,2)$ and $\mathsf{w}=(0,1,0,0)$. Note that the framing is not minuscule, so this $D_{4}$ quiver variety is not treated in [14]. In agreement with our choice of labeling of vertices, we will use $z_{1},z_{2},z_{3}$ (resp. $z_{4}$) for the Kähler variables of $\mathcal{M}^{(1)}$ (resp. $\mathcal{M}^{(2)}$).
@@
$\mathsf{V}_{\mathcal{M}^{(2)}}=\sum_{d_{1},d_{2}\geq 0}\left(\kappa^{2}z_{4}\right)^{d_{1}+d_{2}}\left(\prod_{i,j=2}^{2}\frac{\left(\hbar\frac{a_{i}}{a_{j}}\right)_{d_{i}}}{\left(q\frac{a_{i}}{a_{j}}\right)_{d_{i}}}\right)\left(\prod_{i,j=1}^{2}\frac{\left(q\frac{a_{i}}{a_{j}}\right)_{d_{i}-d_{j}}}{\left(\hbar\frac{a_{i}}{a_{j}}\right)_{d_{i}-d_{j}}}\right)$
see, for example [13]. It is not at all clear from the formula that $V_{\mathcal{M}^{(2)}}$ in independent of $a_{1}$ and $a_{2}$. Nevertheless, it was shown in [13] that
$\mathsf{V}_{\mathcal{M}^{(2)}}=\frac{\Phi\left(\hbar\kappa^{2}z_{4}\right)\Phi\left(\hbar\kappa z_{4}\right)}{\Phi\left(\kappa^{2}z_{4}\right)\Phi\left(\kappa z_{4}\right)}$
in agreement with Conjecture 1.3. This is equivalent to the statement that the Macdonald-Ruijsenaars operators of row type act diagonally on the Macdonald polynomial for the empty partition with a certain eigenvalue, see [34].
@@
$\mathsf{V}_{\mathcal{M}}=\frac{\Phi\left(\hbar\kappa^{2}z_{2}z_{4}\right)\Phi(\hbar\kappa z_{1}z_{2}z_{4})\Phi(\hbar\kappa^{3}z_{2}z_{3}z_{4})\Phi(\hbar\kappa^{2}z_{1}z_{2}z_{3}z_{4})}{\Phi(\kappa^{2}z_{2}z_{4})\Phi(z_{1}z_{2}z_{4}q/\hbar)\Phi(\kappa^{3}z_{2}z_{3}z_{4})\Phi(\kappa^{2}z_{1}z_{2}z_{3}z_{4})}\cdot\frac{\Phi\left(\hbar\kappa^{2}z_{4}\right)\Phi\left(\hbar\kappa z_{4}\right)}{\Phi\left(\kappa^{2}z_{4}\right)\Phi\left(\kappa z_{4}\right)}$
Associated to $\mathcal{M}$ are the weights $\lambda=\varpi_{2}$ and $\mu=\varpi_{2}-2\varpi_{4}=-\epsilon_{3}-\epsilon_{4}$. Using the explicit construction of the $D_{4}$ root system, it is easy to calculate that
$(\alpha,\mu)=\begin{cases}-2&\alpha=\alpha_{4}\\ -1&\alpha\in\{\alpha_{2}+\alpha_{4},\alpha_{1}+\alpha_{2}+\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\}\\ \geq 0&\text{otherwise}\end{cases}$
for any positive root $\alpha$. So we again confirm Conjecture 1.3.
@@
The corresponding Kac-Moody algebra is of indefinite type and is not hyperbolic, see [21], exercise 4.2. We will calculate the vertex function of $\mathcal{M}$ using Corollary 3.5.
@@
and $\kappa=q/\hbar$. This formula was proven in [14]. Using $z_{8}$ (resp. $z_{9}$) as the Kähler parameter of $\mathcal{M}^{(2)}$ (resp. $\mathcal{M}^{(3)}$), we have
@@
In the Kac-Moody algebra, the associated weights are $\lambda=\varpi_{7}$ and
$\mu=-\varpi_{1}+\varpi_{3}+\varpi_{6}+2\varpi_{7}-2\varpi_{8}-2\varpi_{9}$
We can observe directly that all roots $\alpha$ associated to the monomials of $S$ satisfy $(\alpha,\mu)<0$ and with the appropriate magnitude. Using Lemma 6.9, it is a straightforward exercise to check that these are the only positive, real roots satisfying this property.
@@ 4.2 Branching of the nonstationary Ruijsenaars function
Recall from [5] or [13] that there is a bijection
@@
Up to changes in notation, this is exactly equation (3.11) from [35]. In fact, the similarity of (3.11) of [35] with the vertex function of the cotangent bundle of Grassmannian was part of this inspiration for this work.
@@ 5.1 Higgs branches
One important feature of quiver gauge theories (i.e. those $(G,{\bf{N}})$ that arise from a choice of a quiver $Q$ as above) is that we can consider the GIT quotient $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ depending on a stability parameter $\theta$. For generic $\theta$ (see Proposition 2.2) the variety $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ is smooth, symplectic, and is a resolution of singularities of the image of $\pi\colon\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})\rightarrow\mathcal{M}_{H}(G,{\bf{N}})$. So, whenever the map $\pi$ is surjective (see [11] for the criteria), we can extract the Higgs branch $\mathcal{M}_{H}(G,{\bf{N}})$ from the resolution $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$. However, in many cases (for example, if ${\mathsf{w}}=0$), $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ is empty but $\mathcal{M}_{H}(G,{\bf{N}})$ is highly nontrivial. We will denote $\mathcal{M}_{H}(G,{\bf{N}})$ by $\mathcal{M}_{H}$ whenever $(G,{\bf{N}})$ are clear.
@@ 5.2 Coulomb branches
In [8], Braverman, Finkelberg, and Nakajima proposed a mathematical definition of the Coulomb branch $\mathcal{M}_{C}(G,{\bf{N}})$ associated to a gauge theory corresponding to a pair $(G,{\bf{N}})$ as above. The variety $\mathcal{M}_{C}(G,{\bf{N}})$ is an affine Poisson variety. Even for nice quiver gauge theories (e.g., when $Q$ is ADE and $\mathsf{v},\mathsf{w}$ are such that $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ is nonempty), in contrast to $\mathcal{M}_{H}(G,{\bf{N}})$, the varieties $\mathcal{M}_{C}(G,{\bf{N}})$ do not admit symplectic resolutions.
On the other hand, it is known that $\mathcal{M}_{C}(G,{\bf{N}})$ always has symplectic singularities, see [2]. In particular, $\mathcal{M}_{C}(G,{\bf{N}})$ always admits a $\mathbb{Q}$-factorial terminalization $\widetilde{\mathcal{M}}_{C}(G,{\bf{N}})$, which is a partial resolution of $\mathcal{M}_{C}(G,{\bf{N}})$ with nice properties (see [28, Section 4] for the general discussion). We will denote $\mathcal{M}_{C}(G,{\bf{N}})$ by $\mathcal{M}_{C}$ whenever $(G,{\bf{N}})$ are clear.
@@
The group $G_{\mathcal{O}}$ acts naturally on the space $\mathcal{R}$. In [8, 2(ii)], the authors define the equivariant Borel-Moore homology $H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ of $\mathcal{R}$. Moreover, in [8, Section 3], they endow $H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ with the algebra structure given by convolution *. Finally, they prove that * is commutative and define:

### v2

@@ 3.5 From branching to factorization
Due to form of the obstruction theory, the $q=\hbar$ specialization of vertex functions with descendant $1$ depends on $\hbar$ and not on any other equivariant parameters.
@@
Another interesting special case occurs when $\mathsf{w}^{(2)}_{\star_{2}}=1$ and $\mathsf{w}^{(2)}_{i}=0$ otherwise, which most closely resembles the original slant sum constructions of [undefav].
@@ 4.1 Factorization in zero-dimensional case
In this section, we will follow standard notation and denote simple (co)roots by $\alpha_{i}$ $(\alpha_{i}^{\vee}$) and fundamental (co)weights by $\varpi_{i}$ ($\varpi_{i}^{\vee})$. We will also denote $\kappa=q/\hbar$.
Corollary 3.5 provides a strategy to approach Conjecture 1.5. This conjecture is known only for type $A$ quiver varieties, proven in specific cases in [undefo] and [undefs], and later proven for all type $A$ in [undefe], and type $D$ quiver varieties with framings only at minuscule vertices in [undefp].
@@
$\mathsf{V}_{\mathcal{M}}(z)$ $=\sum_{d_{i,j}}\left(\kappa\right)^{N(d)}z_{1}^{d_{1,1}}z_{2}^{d_{2,1}+d_{2,2}}z_{3}^{d_{3,1}}\frac{\left(\hbar\right)_{d_{2,1}}\left(\hbar^{2}\right)_{d_{2,2}}}{\left(q\right)_{d_{2,1}}\left(q\hbar\right)_{d_{2,2}}}\cdot$
@@
$\mathsf{V}_{\mathcal{M}}(z)=\frac{\Phi\left(\hbar\kappa z_{2}\right)\Phi(\hbar z_{1}z_{2})\Phi(\hbar\kappa^{2}z_{2}z_{3})\Phi(\hbar\kappa z_{1}z_{2}z_{3})}{\Phi(\kappa z_{2})\Phi(z_{1}z_{2})\Phi(\kappa^{2}z_{2}z_{3})\Phi(\kappa z_{1}z_{2}z_{3})}$
The coweights $\lambda$ and $\mu$ in this case are $\lambda=\varpi_{2}^{\vee}$ and $\mu=-\varpi_{2}^{\vee}$. It is easy to see that
$\langle\alpha,\mu\rangle=\begin{cases}1&\text{if $\alpha\in\{-\alpha_{2},-\alpha_{1}-\alpha_{2},\alpha_{2}-\alpha_{3},\alpha_{1}-\alpha_{2}-\alpha_{3}\}$}\\ \geq 0&\text{otherwise}\end{cases}$
for any negative root $\alpha$. Recalling that $e^{-\alpha_{1}}=\kappa^{-1}z_{1}^{-1}$, $e^{-\alpha_{2}}=\kappa z_{2}^{-1}$, and $e^{-\alpha_{3}}=\kappa z_{3}^{-1}$, and that the vertex $\mathsf{V}_{\mathcal{M}}$ is evaluated at $z^{-1}$, this confirms Conjecture 1.5 in this case.
Suppose that $\mathcal{M}^{(1)}$ and $\mathcal{M}^{(2)}$ are zero-dimensional quiver varieties whose vertex functions are known to factorize into a product of $q$-binomial functions. By the discussion above, they could be any type $A$ quiver varieties or type $D$ quiver varieties with framings only at minuscule vertices. Suppose that $\star_{1}$ and $\star_{2}$ are compatible vertices, that $\mathcal{M}^{(1)}$ is split over $\star_{1}$, and that $\theta^{(2)}=\pm(1,1,\ldots,1)$. Consider $\mathcal{M}:=\mathcal{M}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathcal{M}^{(2)}$.
With these assumptions, Corollary 3.5 implies the following, see also Remark 3.6.
@@
Let $\mathcal{M}^{(1)}$ be the type $A_{3}$ quiver variety with $\mathsf{v}^{(1)}=(1,2,1)$ and $\mathsf{w}^{(1)}=(0,1,0)$. Let $\mathcal{M}^{(2)}$ be the type $A_{1}$ quiver variety with $\mathsf{v}^{(2)}=(2)$ and $\mathsf{w}^{(2)}=(2)$. Let $\star_{1}$ be the second vertex and let $\star_{2}$ be the first (and only) vertex. As a $\mathsf{T}^{(1)}$-representation, $\mathcal{V}_{\star_{1}}=1+\hbar$, so $\mathcal{M}^{(1)}$ is split over $\star_{1}$.
Let $\mathcal{M}=\mathcal{M}^{(1)}{}_{\star_{1}}\#_{\star_{2}}\mathcal{M}^{(2)}$. It is a $D_{4}$ quiver variety with $\mathsf{v}=(1,2,1,2)$ and $\mathsf{w}=(0,1,0,0)$. Note that the framing is not minuscule, so this $D_{4}$ quiver variety is not treated in [undefp]. In agreement with our choice of labeling of vertices, we will use $z_{1},z_{2},z_{3}$ (resp. $z_{4}$) for the Kähler variables of $\mathcal{M}^{(1)}$ (resp. $\mathcal{M}^{(2)}$).
@@
$\mathsf{V}_{\mathcal{M}^{(2)}}(z)=\sum_{d_{1},d_{2}\geq 0}\left(\kappa^{2}z_{4}\right)^{d_{1}+d_{2}}\left(\prod_{i,j=2}^{2}\frac{\left(\hbar\frac{a_{i}}{a_{j}}\right)_{d_{i}}}{\left(q\frac{a_{i}}{a_{j}}\right)_{d_{i}}}\right)\left(\prod_{i,j=1}^{2}\frac{\left(q\frac{a_{i}}{a_{j}}\right)_{d_{i}-d_{j}}}{\left(\hbar\frac{a_{i}}{a_{j}}\right)_{d_{i}-d_{j}}}\right)$
see, for example [undefo]. It is not at all clear from the formula that $V_{\mathcal{M}^{(2)}}$ in independent of $a_{1}$ and $a_{2}$. Nevertheless, it was shown in [undefo] that
$\mathsf{V}_{\mathcal{M}^{(2)}}(z)=\frac{\Phi\left(\hbar\kappa^{2}z_{4}\right)\Phi\left(\hbar\kappa z_{4}\right)}{\Phi\left(\kappa^{2}z_{4}\right)\Phi\left(\kappa z_{4}\right)}$
in agreement with Conjecture 1.5. This is equivalent to the statement that the Macdonald-Ruijsenaars operators of row type act diagonally on the Macdonald polynomial for the empty partition with a certain eigenvalue, see [undefar].
@@
$\mathsf{V}_{\mathcal{M}}(z)=\frac{\Phi\left(\hbar\kappa^{2}z_{2}z_{4}\right)\Phi(\hbar\kappa z_{1}z_{2}z_{4})\Phi(\hbar\kappa^{3}z_{2}z_{3}z_{4})\Phi(\hbar\kappa^{2}z_{1}z_{2}z_{3}z_{4})}{\Phi(\kappa^{2}z_{2}z_{4})\Phi(z_{1}z_{2}z_{4}q/\hbar)\Phi(\kappa^{3}z_{2}z_{3}z_{4})\Phi(\kappa^{2}z_{1}z_{2}z_{3}z_{4})}\cdot\frac{\Phi\left(\hbar\kappa^{2}z_{4}\right)\Phi\left(\hbar\kappa z_{4}\right)}{\Phi\left(\kappa^{2}z_{4}\right)\Phi\left(\kappa z_{4}\right)}$
Associated to $\mathcal{M}$ are the coweights $\lambda=\varpi_{2}^{\vee}$ and $\mu=\varpi_{2}^{\vee}-2\varpi_{4}^{\vee}$. Using the explicit construction of the $D_{4}$ root system, it is easy to calculate that
$\langle\alpha,\mu\rangle=\begin{cases}2&\alpha=-\alpha_{4}\\ 1&\alpha\in\{-\alpha_{2}-\alpha_{4},-\alpha_{1}-\alpha_{2}-\alpha_{4},-\alpha_{2}-\alpha_{3}-\alpha_{4},-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}\}\\ \leq 0&\text{otherwise}\end{cases}$
for any negative root $\alpha$. So we again confirm Conjecture 1.5.
@@
The corresponding Kac-Moody algebra is of indefinite type and is not hyperbolic, see [undefaa], exercise 4.2. We will calculate the vertex function of $\mathcal{M}$ using Corollary 3.5.
@@
and $\kappa=q/\hbar$. This formula was proven in [undefp]. Using $z_{8}$ (resp. $z_{9}$) as the Kähler parameter of $\mathcal{M}^{(2)}$ (resp. $\mathcal{M}^{(3)}$), we have
@@
In the Kac-Moody algebra, the associated coweights are $\lambda=\varpi_{7}^{\vee}$ and
$\mu=-\varpi_{1}^{\vee}+\varpi_{3}^{\vee}+\varpi_{6}^{\vee}+2\varpi_{7}^{\vee}-2\varpi_{8}^{\vee}-2\varpi_{9}^{\vee}$
We can observe directly that all roots $\alpha$ associated to the monomials of $S$ satisfy $\langle\alpha,\mu\rangle>0$ and with the appropriate magnitude. Using Lemma 7.10, it is a straightforward exercise to check that these are the only negative, real roots satisfying this property.
@@ 4.2 Branching of the nonstationary Ruijsenaars function
Recall from [undefe] or [undefo] that there is a bijection
@@
Up to changes in notation, this is exactly equation (3.11) from [undefas]. In fact, the similarity of (3.11) of [undefas] with the vertex function of the cotangent bundle of Grassmannian was part of this inspiration for this work.
@@ 5.1 Higgs branches
One important feature of quiver gauge theories (i.e. those $(G,{\bf{N}})$ that arise from a choice of a quiver $Q$ as above) is that we can consider the GIT quotient $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ depending on a stability parameter $\theta$. For generic $\theta$ (see Proposition 2.2) the variety $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ is smooth, symplectic, and is a resolution of singularities of the image of $\pi\colon\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})\rightarrow\mathcal{M}_{H}(G,{\bf{N}})$. So, whenever the map $\pi$ is surjective (see [undefm] for the criteria), we can extract the Higgs branch $\mathcal{M}_{H}(G,{\bf{N}})$ from the resolution $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$. However, in many cases (for example, if ${\mathsf{w}}=0$), $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ is empty but $\mathcal{M}_{H}(G,{\bf{N}})$ is highly nontrivial. We will denote $\mathcal{M}_{H}(G,{\bf{N}})$ by $\mathcal{M}_{H}$ whenever $(G,{\bf{N}})$ are clear.
@@ 5.2 Coulomb branches
In [undefj], Braverman, Finkelberg, and Nakajima proposed a mathematical definition of the Coulomb branch $\mathcal{M}_{C}(G,{\bf{N}})$ associated to a gauge theory corresponding to a pair $(G,{\bf{N}})$ as above. The variety $\mathcal{M}_{C}(G,{\bf{N}})$ is an affine Poisson variety. Even for nice quiver gauge theories (e.g., when $Q$ is ADE and $\mathsf{v},\mathsf{w}$ are such that $\mathcal{M}_{Q,\theta}(\mathsf{v},\mathsf{w})$ is nonempty), in contrast to $\mathcal{M}_{H}(G,{\bf{N}})$, the varieties $\mathcal{M}_{C}(G,{\bf{N}})$ do not admit symplectic resolutions.
On the other hand, it is known that $\mathcal{M}_{C}(G,{\bf{N}})$ always has symplectic singularities, see [undefa]. We will denote $\mathcal{M}_{C}(G,{\bf{N}})$ by $\mathcal{M}_{C}$ whenever $(G,{\bf{N}})$ are clear.
@@
The group $G_{\mathcal{O}}$ acts naturally on the space $\mathcal{R}$. In [undefj, 2(ii)], the authors define the equivariant Borel-Moore homology $H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ of $\mathcal{R}$. Moreover, in [undefj, Section 3], they endow $H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ with the algebra structure given by convolution *. Finally, they prove that * is commutative and define: